# Batalin-Fradkin-Vilkovisky Quantization of FLPR model

### v1

@@ 5 Conclusions
We have performed the BFV quantization of the FLPR model in both polar and Cartesian coordinates using admissible gauge conditions. The FLPR model is endowed with first-class constraints à la Dirac’s prescription for classification of constraints. In order to apply BFV formalism, we have introduced a pair of ghost and anti-ghost variables along with auxiliary variables. The conserved and nilpotent charge has been constructed using constraints. Whereas the fermionic gauge-fixing function is being put together by means of admissible gauge-fixing conditions. We have obtained the effective action, in the extended phase space, by the usual Legendre transformation. By eliminating a few variables using Gaussian integration, we have established the BRST invariant gauge-fixed effective action and corresponding BRST symmetries of the FLPR model in both polar and Cartesian coordinates. We have also indicated that the effective action has a discrete symmetry in the ghost sector. In addition, using the physicality criterion, we have shown that the physical states of the theory are annihilated by the first-class constraints, which is consistent with the Dirac formalism.

### v2

@@ 7 Conclusions
We have performed the BFV quantization of the FLPR model in both polar and Cartesian coordinates using admissible gauge conditions. The FLPR model is endowed with first-class constraints à la Dirac’s prescription for classification of constraints. In order to apply the BFV formalism, we have introduced a pair of ghost and anti-ghost variables along with auxiliary variables. The conserved and nilpotent charge has been constructed using constraints. Whereas the fermionic gauge-fixing function is being put together by means of admissible gauge-fixing conditions. We have obtained the effective action, in the extended phase space, by the usual Legendre transformation. By eliminating a few variables using Gaussian integration, we have established the BRST invariant gauge-fixed effective action and corresponding BRST symmetries of the FLPR model in both polar and Cartesian coordinates. We have also indicated that the effective action has a discrete symmetry in the ghost sector. Using the physicality criterion, we have shown that the physical states of the theory are annihilated by the first-class constraints, which is consistent with the Dirac formalism.
Further, we have generalized the infinitesimal BRST symmetries by making the transformation parameter finite and field dependent, thereby establishing the FFBRST symmetries of the FLPR model. Although the action is invariant under such transformations, the field dependence of the finite parameter breaks the invariance of generating functional. We have chosen an appropriate FFBRST parameter and computed the non-trivial Jacobian of the FFBRST transformation. Thus, we are able to show that with a judicious choice of the finite field-dependent parameter, FFBRST transformations can bridge between generating functional corresponding to two different actions. We have demonstrated the interconnection between the BFV-BRST gauge-fixed effective action and the classical gauge invariant action within FFBRST formulation.